OpenStudy (anonymous):

Using complete sentences, describe the characteristics and parts of sigma notation. Please also provide a unique example.

OpenStudy (anonymous):

you mean like Sum?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Sigma, after all is just a Greek letter "s"

OpenStudy (anonymous):

Unless Sigma as referring to Sum

OpenStudy (anonymous):

its referring to like summation notation

OpenStudy (anonymous):

\[\sum _{n=0}^N a_n\]

OpenStudy (anonymous):

In the above Summation 1) begin at n=0 2) ends at N 3) summation of a n

OpenStudy (anonymous):

\[\sum _{n=0}^5 n=0+1+2+3+4+5\]